# MMGT: Motion Mask Guided Two-Stage Network for Co-Speech Gesture Video Generation

### v1

@@ IV-E Ethical Considerations and Mitigation Strategies
This study identifies two key societal risks in audio-driven portrait animation technology: potential misuse for creating deceptive deepfakes and privacy violations through unauthorized use of personal biometric data. Addressing these concerns requires establishing ethical guidelines, ensuring transparent deployment, securing explicit user consent, and implementing robust data governance. These measures aim to align technological development with social accountability while promoting socially beneficial applications.
V conclusions.
We propose a Motion Mask-Guided Two-Stage Network (MMGT) for Co-Speech Gesture Video Generation, integrating motion masks and features extracted from audio. The Spatial Mask-Guided Audio Pose Generation (SMGA) network captures large-scale facial and gesture movements, while the Motion Masked Hierarchical Audio Attention (MM-HAA) mechanism refines fine-grained details within a stabilized diffusion framework. Experiments show that MMGT produces natural, high-quality videos with enhanced motion realism and texture accuracy.

### v2

@@ IV-E Ethical Considerations and Mitigation Strategies
This study highlights two major societal risks in audio-driven portrait animation technology: potential misuse for creating deceptive deepfakes and privacy breaches through unauthorized use of personal biometric data. Addressing these issues requires establishing ethical standards, ensuring transparent implementation, obtaining clear user consent, and enforcing strong data governance. These steps aim to align technological progress with social responsibility while encouraging socially beneficial uses.
Figure 15: Qualitative Comparison of Collaborative Speech Gesture Video Generation Methods in the Face of Complex Interactions.
V Discussion and limitations
Although MMGT significantly improves overall motion realism and texture fidelity, it still has shortcomings when handling fine physical interactions (such as hand-to-face contact or hand-to-object manipulation). As shown in Fig. 15, we found that this phenomenon is common in existing methods[20, 54]. We believe that this defect may mainly stem from the lack of dense annotated training data for such events and the lack of explicit contact constraints in the current architecture.
VI Conclusions.
We propose a Motion Mask-Guided Two-Stage Network (MMGT) for Co-Speech Gesture Video Generation, which combines motion masks with features extracted from audio. The Spatial Mask-Guided Audio2Pose Generation (SMGA) network captures large-scale facial and hand gesture movements, while the Motion Masked Hierarchical Audio Attention (MM-HAA) mechanism refines fine details within a stabilized diffusion framework. Experiments demonstrate that MMGT produces natural, high-quality videos with improved motion realism and texture accuracy.